OpenStudy (anonymous):

What is 4 whole times 1/8 ? Can you explain how to do it also?

myininaya (myininaya):

\[4 \cdot \frac{1}{8}=\frac{4}{8}=\frac{4}{2 \cdot 4}=\frac{\not{4}}{2 \not{4}}=\frac{1}{2}\]

OpenStudy (anonymous):

Oh I see. Thanks